Consider the following differential equation system which describes the interaction between two species.


 Ẋ = 0.2x βˆ’ 0.0002x2 βˆ’ 0.006π‘₯𝑦

𝑦̇ = 0.015𝑦 + 0.0008π‘₯𝑦


a) What type of interaction is modeled by the system? What is the significance of the variables?

b) Find the equilibrium points and analyze their stability


1
Expert's answer
2021-05-23T16:43:23-0400

The type of behavior is competing for exclusion. Variables have the value of quantitative values of two independent factors.


"\frac{\delta X}{\delta x}=0.2-0.0004x-0.006y=0"

"\frac{\delta Y}{\delta y}=0.015+0.0008y=0"

"x=18.75"

"y=32"